Solve each system of equations. If no solution explain why!
Substitution method
x-y=3
2x-3y=-1
*****************
Addition Method
3x+2y=1
5x+3y=3
*****************
Any method
1/4a - 3/4b = -1
1/3a + b = 5/3
*****************
Any method
3x+2y=0
2x-y+3z=8
5x+3y+z=4
*****************
Solve systems of linear inequalities by graphing
3x+y>8
x-2y>5
******************
any method
5x-3y=3
7x+y=25

Answer by Alan3354(69443)

Solve each system of equations. If no solution explain why!
Substitution method
x-y=3
2x-3y=-1
------------
x-y=3 --> x = y+3
Sub for x in the 2nd eqn
*****************
Addition Method
3x+2y=1
5x+3y=3
Multiply eqn 1 by 3, eqn 2 by 2, then subtract
--------------
*****************
Any method
1/4a - 3/4b = -1
1/3a + b = 5/3
Multiply eqn 1 by 4 and eqn 2 by 3, then add
*****************
Any method
3x+2y=0
2x-y+3z=8
5x+3y+z=4
I would use determinants.
To check you answers: (-2,3,5)
*****************
Solve systems of linear inequalities by graphing
3x+y>8
x-2y>5
Graph them.
******************
any method
5x-3y=3
7x+y=25
I would use elimination, as above.